Let x=18
13(x) + 18y=60

Answer by colliefan(242)

13(x) + 18y=60
13(18)+18y = 60
13(18)+18y = 60
234+18y=60
234+18y-234=60-234
18y+0=-174
18y=-174
18y*1/18=-174*1/18
y*1=-174/18
y=-174/18
y=-87/9